OpenStudy (anonymous):

Flaws in a carpet tend to occur randomly and independently at a rate of one every 300 square feet. What is the probability that a carpet that is 7 feet by 11 feet contains no flaws?

OpenStudy (anonymous):

Well according to the math there would be no flaws

OpenStudy (kropot72):

Using the Poisson distribution the required probability is \[P(X=0)=\frac{e ^{-\frac{77}{300}}(\frac{77}{300})^{0}}{0!}=0.77336\]